The GNU Binary Utilities

Version 2.2 May 1993

Roland H. Pesch Cygnus Support

Cygnus Support Revision: 1.22 TEXinfo 2.107

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1991, 1992, 1993 Free Software Foundation, Inc.

Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies.

Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided also that the entire resulting derived work is distributed under the terms of a permission notice identical to this one.

Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions.

1 Introduction

This brief manual contains preliminary documentation for the GNU binary utilities (collectively version 2.2):

ar Create, modify, and extract from archives objcopy Copy and translate object files nm List symbols from object files objdump Display information from object files ranlib Generate index to archive contents size List section sizes and total size strip Discard symbols 2 GNU Binary Utilities

2 ar

ar [-]pmod [ membername ] archive file…

     ar -M [ <mri-script ]

The GNU ar program creates, modifies, and extracts from archives. An archive is a single file holding a collection of other files in a structure that makes it possible to retrieve the original individual files (called members of the archive).

The original files’ contents, mode (permissions), timestamp, owner, and group are preserved in the archive, and can be restored on extraction.

GNU ar can maintain archives whose members have names of any length; however, depending on how ar is configured on your system, a limit on member-name length may be imposed for compatibility with archive formats maintained with other tools. If it exists, the limit is often 15 characters (typical of formats related to a.out) or 16 characters (typical of formats related to coff).

ar is considered a binary utility because archives of this sort are most often used as libraries holding commonly needed subroutines.

ar creates an index to the symbols defined in relocatable object modules in the archive when you specify the modifier ‘s’. Once created, this index is updated in the archive whenever ar makes a change to its contents (save for the ‘q’ update operation). An archive with such an index speeds up linking to the library, and allows routines in the library to call each other without regard to their placement in the archive.

You may use ‘nm -s’ or ‘nm --print-armap’ to list this index table. If an archive lacks the table, another form of ar called ranlib can be used to add just the table.

GNU ar is designed to be compatible with two different facilities. You can control its activity using command-line options, like the different varieties of ar on Unix systems; or, if you specify the single command-line option ‘-M’, you can control it with a script supplied via standard input, like the MRI “librarian” program.

2.1 Controlling ar on the command line

ar [-]pmod [ membername ] archive file…

When you use ar in the Unix style, ar insists on at least two arguments to execute: one keyletter specifying the operation (optionally accompanied by other keyletters specifying modifiers), and the archive name to act on.

Most operations can also accept further file arguments, specifying particular files to operate on.

GNU ar allows you to mix the operation code p and modifier flags mod in any order, within the first command-line argument.

If you wish, you may begin the first command-line argument with a dash.

The p keyletter specifies what operation to execute; it may be any of the following, but you must specify only one of them:

d Delete modules from the archive. Specify the names of modules to be deleted as file…; the archive is untouched if you specify no files to delete. If you specify the ‘v’ modifier, ar lists each module as it is deleted.

m Use this operation to move members in an archive.

The ordering of members in an archive can make a difference in how programs are linked using the library, if a symbol is defined in more than one member. If no modifiers are used with m, any members you name in the file arguments are moved to the end of the archive; you can use the ‘a’, ‘b’, or ‘i’ modifiers to move them to a specified place instead.

p Print the specified members of the archive, to the standard output file. If the ‘v’ modifier is specified, show the member name before copying its contents to standard output. If you specify no file arguments, all the files in the archive are printed.

q Quick append; add the files file… to the end of archive, without checking for replacement. The modifiers ‘a’, ‘b’, and ‘i’ do not affect this operation; new members are always placed at the end of the archive. The modifier ‘v’ makes ar list each file as it is appended.

Since the point of this operation is speed, the archive’s symbol table index is not updated, even if it already existed; you can use ‘ar s’ or ranlib explicitly to update the symbol table index.

r Insert the files file… into archive (with replacement). This operation differs from ‘q’ in that any previously existing members are deleted if their names match those being added. If one of the files named in file… doesn’t exist, ar displays an error message, and leaves undisturbed any existing members of the archive matching that name. By default, new members are added at the end of the file; but you may use one of the modifiers ‘a’, ‘b’, or ‘i’ to request placement relative to some existing member. The modifier ‘v’ used with this operation elicits a line of output for each file inserted, along with one of the letters ‘a’ or ‘r’ to indicate whether the file was appended (no old member deleted) or replaced.

t Display a table listing the contents of archive, or those of the files listed in file… that are present in the archive. Normally only the member name is shown; if you also want to see the modes (permissions), timestamp, owner, group, and size, you can request that by also specifying the ‘v’ modifier. If you do not specify a file, all files in the archive are listed. If there is more than one file with the same name (say, ‘fie’) in an archive (say ‘b.a’), ‘ar t b.a fie’ lists only the first instance; to see them all, you must ask for a complete listing—in our example, ‘ar t b.a’.

x Extract members (named file) from the archive. You can use the ‘v’ modifier with this operation, to request that ar list each name as it extracts it. If you do not specify a file, all files in the archive are extracted.

A number of modifiers (mod) may immediately follow the p keyletter, to specify variations on an operation’s behavior:

a Add new files after an existing member of the archive. If you use the modifier ‘a’, the name of an existing archive member must be present as the membername argument, before the archive specification.

b Add new files before an existing member of the archive. If you use the modifier ‘b’, the name of an existing archive member must be present as the membername argument, before the archive specification. (same as ‘i’).

c Create the archive. The specified archive is always created if it didn’t exist, when you request an update. But a warning is issued unless you specify in advance that you expect to create it, by using this modifier. i Insert new files before an existing member of the archive. If you use the modifier ‘i’, the

name of an existing archive member must be present as the membername argument, before the archive specification. (same as ‘b’).

l This modifier is accepted but not used.

o Preserve the original dates of members when extracting them. If you do not specify this modifier, files extracted from the archive are stamped with the time of extraction.

s Write an object-file index into the archive, or update an existing one, even if no other change is made to the archive. You may use this modifier flag either with any operation, or alone. Running ‘ar s’ on an archive is equivalent to running ‘ranlib’ on it.

u Normally, ‘ar r’… inserts all files listed into the archive. If you would like to insert only those of the files you list that are newer than existing members of the same names, use this modifier. The ‘u’ modifier is allowed only for the operation ‘r’ (replace). In particular, the combination ‘qu’ is not allowed, since checking the timestamps would lose any speed advantage from the operation ‘q’.

v This modifier requests the verbose version of an operation. Many operations display additional information, such as filenames processed, when the modifier ‘v’ is appended.

V This modifier shows the version number of ar.

2.2 Controlling ar with a script

     ar -M [ <script ]

If you use the single command-line option ‘-M’ with ar, you can control its operation with a rudimentary command language. This form of ar operates interactively if standard input is coming directly from a terminal. During interactive use, ar prompts for input (the prompt is ‘AR >’), and continues executing even after errors. If you redirect standard input to a script file, no prompts are issued, and ar abandons execution (with a nonzero exit code) on any error.

The ar command language is not designed to be equivalent to the command-line options; in fact, it provides somewhat less control over archives. The only purpose of the command language is to ease the transition to GNU ar for developers who already have scripts written for the MRI “librarian” program.

The syntax for the ar command language is straightforward:

• commands are recognized in upper or lower case; for example, LIST is the same as list. In the following descriptions, commands are shown in upper case for clarity.

• a single command may appear on each line; it is the first word on the line.

• empty lines are allowed, and have no effect.

• comments are allowed; text after either of the characters ‘*’ or ‘;’ is ignored.

• Whenever you use a list of names as part of the argument to an ar command, you can separate the individual names with either commas or blanks. Commas are shown in the explanations below, for clarity.

• ‘+’ is used as a line continuation character; if ‘+’ appears at the end of a line, the text on the following line is considered part of the current command.

Here are the commands you can use in ar scripts, or when using ar interactively. Three of them have special significance:

OPEN or CREATE specify a current archive, which is a temporary file required for most of the other commands.

SAVE commits the changes so far specified by the script. Prior to SAVE, commands affect only the temporary copy of the current archive.

ADDLIB archive ADDLIB archive (module, module, … module)

Add all the contents of archive (or, if specified, each named module from archive) to the current archive. Requires prior use of OPEN or CREATE.

ADDMOD file, file, … file Add each named file as a module in the current archive. Requires prior use of OPEN or CREATE.

CLEAR Discard the contents of the current archive, cancelling the effect of any operations since

the last SAVE. May be executed (with no effect) even if no current archive is specified.

CREATE archive Creates an archive, and makes it the current archive (required for many other commands). The new archive is created with a temporary name; it is not actually saved as archive until you use SAVE. You can overwrite existing archives; similarly, the contents of any existing file named archive will not be destroyed until SAVE.

DELETE module, module, … module Delete each listed module from the current archive; equivalent to ‘ar -d archive module … module’. Requires prior use of OPEN or CREATE.

DIRECTORY archive (module, … module) DIRECTORY archive (module, … module) outputfile

List each named module present in archive. The separate command VERBOSE specifies the form of the output: when verbose output is off, output is like that of ‘ar -t archive module…’. When verbose output is on, the listing is like ‘ar -tv archive module…’. Output normally goes to the standard output stream; however, if you specify outputfile as a final argument, ar directs the output to that file.

END Exit from ar, with a 0 exit code to indicate successful completion. This command does not save the output file; if you have changed the current archive since the last SAVE command, those changes are lost.

EXTRACT module, module, … module Extract each named module from the current archive, writing them into the current directory as separate files. Equivalent to ‘ar -x archive module…’. Requires prior use of OPEN or CREATE.

LIST Display full contents of the current archive, in “verbose” style regardless of the state of VERBOSE. The effect is like ‘ar tv archive’). (This single command is a GNU ld enhancement, rather than present for MRI compatibility.) Requires prior use of OPEN or CREATE.

OPEN archive Opens an existing archive for use as the current archive (required for many other commands). Any changes as the result of subsequent commands will not actually affect archive until you next use SAVE.

REPLACE module, module, … module In the current archive, replace each existing module (named in the REPLACE arguments) from files in the current working directory. To execute this command without errors, both the file, and the module in the current archive, must exist. Requires prior use of OPEN or CREATE.

VERBOSE Toggle an internal flag governing the output from DIRECTORY. When the flag is on,

DIRECTORY output matches output from ‘ar -tv ’. . ..

SAVE Commit your changes to the current archive, and actually save it as a file with the name specified in the last CREATE or OPEN command. Requires prior use of OPEN or CREATE.

3 objcopy

     objcopy [ -F format | --format=format ]
             [ -I format | --input-format=format ]
             [ -O format | --output-format=format ]
             [ -S | --strip-all ]  [ -g | --strip-debug ]
             [ -x | --discard-all ]  [ -X | --discard-locals ]
             [ -v | --verbose ]  [ -V | --version ]

infile [outfile]

The GNU objcopy utility copies the contents of an object file to another. objcopy uses the GNU BFD Library to read and write the object files. It can write the destination object file in a format different from that of the source object file. The exact behavior of objcopy is controlled by command-line options.

objcopy creates temporary files to do its translations and deletes them afterward. objcopy uses BFD to do all its translation work; it knows about all the formats BFD knows about, and thus is able to recognize most formats without being told explicitly. See section “BFD” in Using LD.

infile outfile The source and output files respectively. If you do not specify outfile, objcopy creates

a temporary file and destructively renames the result with the name of the input file.

-I format

--input-format=format

Consider the source file’s object format to be format, rather than attempting to deduce it.

-O format

--output-format=format

Write the output file using the object format format.

-F format

--format=format

Use format as the object format for both the input and the output file; i.e. simply transfer data from source to destination with no translation.

-S
--strip-all

Do not copy relocation and symbol information from the source file.

-g
--strip-debug

Do not copy debugging symbols from the source file.

-x
--discard-all

Do not copy non-global symbols from the source file.

-X
--discard-locals

Do not copy compiler-generated local symbols. (These usually start with ‘L’ or ‘.’.)

-V
--version

Show the version number of objcopy.

-v
--verbose

Verbose output: list all object files modified. In the case of archives, ‘objcopy -V’ lists all members of the archive.

Chapter 4: ld 11

4 ld

The GNU linker ld is now described in a separate manual. See section “Overview” in Using LD: the GNU linker.

12 GNU Binary Utilities

5 nm

     nm [ -a | --debug-syms ]  [ -g | --extern-only ]
        [ -s | --print-armap ]  [ -o | --print-file-name ]
        [ -n | --numeric-sort ]  [ -p | --no-sort ]
        [ -r | --reverse-sort ]  [ -u | --undefined-only ]
        [ --target=bfdname ]

[ objfile…]

GNU nm lists the symbols from object files objfile. . ..

The long and short forms of options, shown here as alternatives, are equivalent.

objfile… Object files whose symbols are to be listed. If no object files are listed as arguments,

nm assumes ‘a.out’.

-a
--debug-syms

Display debugger-only symbols; normally these are not listed.

-g
--extern-only

Display only external symbols.

-p
--no-sort

Don’t bother to sort the symbols in any order; just print them in the order encountered.

-n
--numeric-sort

Sort symbols numerically by their addresses, rather than alphabetically by their names.

-s
--print-armap

When listing symbols from archive members, include the index: a mapping (stored in the archive by ar or ranlib) of which modules contain definitions for which names.

-o
--print-file-name

Precede each symbol by the name of the input file where it was found, rather than identifying the input file once only before all of its symbols.

-r
--reverse-sort

Reverse the order of the sort (whether numeric or alphabetic); let the last come first.

--target=bfdname

Specify an object code format other than your system’s default format. See Chapter 6 [objdump], page 15, for information on listing available formats.

-u
--undefined-only

Display only undefined symbols (those external to each object file).

6 objdump

     objdump [ -a ]  [ -b bfdname ]  [ -d ]  [ -f ]
             [ -h | --header ]  [ -i ]  [ -j section ]  [ -l ]
             [ -m machine ]  [ -r | --reloc ]  [ -s ]
             [ --stabs ]  [ -t | --syms ]  [ -x ]

objfile…

objdump displays information about one or more object files. The options control what particular information to display. This information is mostly useful to programmers who are working on the compilation tools, as opposed to programmers who just want their program to compile and work.

The long and short forms of options, shown here as alternatives, are equivalent.

objfile… The object files to be examined. When you specify archives, objdump shows information

on each of the member object files.

-a If any of the objfile files are archives, display the archive header information (in a format similar to ‘ls -l’). Besides the information you could list with ‘ar tv’, ‘objdump -a’ shows the object file format of each archive member.

-b bfdname Specify that the object-code format for the object files is bfdname. This option may not be necessary; objdump can automatically recognize many formats. For example,

               objdump -b oasys -m vax -h fu.o

displays summary information from the section headers (‘-h’) of ‘fu.o’, which is explicitly identified (‘-m’) as a VAX object file in the format produced by Oasys compilers. You can list the formats available with the ‘-i’ option.

-d Disassemble. Display the assembler mnemonics for the machine instructions from obj-

file.

-f File header. Display summary information from the overall header of each of the objfile

files.

-h

--header Header. Display summary information from the section headers of the object file. -i Display a list showing all architectures and object formats available for specification

with ‘-b’ or ‘-m’.

-j name Display information only for section name.

-l Label the display (using debugging information) with the source filename and line

numbers corresponding to the object code shown.

-m machine Specify that the object files objfile are for architecture machine. You can list available architectures using the ‘-i’ option.

-r

--reloc Relocation. Print the relocation entries of the file. -s Display the full contents of any sections requested. --stabs Display the full contents of any sections requested. Display the contents of the .stab

and .stab.index and .stab.excl sections from an ELF file. This is only useful on systems (such as Solaris 2.0) in which .stab debugging symbol-table entries are carried in an ELF section. In most other file formats, debugging symbol-table entries are interleaved with linkage symbols, and are visible in the ‘--syms’ output.

-t

--syms Symbol Table. Print the symbol table entries of the file. This is similar to the infor-

mation provided by the ‘nm’ program.

-x Display all available header information, including the symbol table and relocation entries. Using ‘-x’ is equivalent to specifying all of ‘-a -f -h -r -t’.

7 ranlib

     ranlib [-vV] archive

ranlib generates an index to the contents of an archive and stores it in the archive. The index lists each symbol defined by a member of an archive that is a relocatable object file.

You may use ‘nm -s’ or ‘nm --print-armap’ to list this index.

An archive with such an index speeds up linking to the library and allows routines in the library to call each other without regard to their placement in the archive.

The GNU ranlib program is another form of GNU ar; running ranlib is completely equivalent to executing ‘ar -s’. See Chapter 2 [ar], page 3.

-v

-V Show the version number of ranlib.

18 GNU Binary Utilities

8 size

size [ -A | -B | --format=compatibility ]
[ --help ]  [ -d | -o | -x | --radix=number ]
[ --target=bfdname ]  [ -V | --version ]

objfile…

The GNU size utility lists the section sizes—and the total size—for each of the object or archive files objfile in its argument list. By default, one line of output is generated for each object file or each module in an archive.

The command line options have the following meanings:

objfile… The object files to be examined.

-A
-B

--format=compatibility

Using one of these options, you can choose whether the output from GNU size resembles output from System V size (using ‘-A’, or ‘--format=sysv’), or Berkeley size (using ‘-B’, or ‘--format=berkeley’). The default is the one-line format similar to Berkeley’s.

Here is an example of the Berkeley (default) format of output from size:

size --format Berkeley ranlib size
text    data    bss     dec     hex     filename
294880  81920   11592   388392  5ed28   ranlib
294880  81920   11888   388688  5ee50   size

This is the same data, but displayed closer to System V conventions:

size --format SysV ranlib size
ranlib  :
section         size         addr
.text         294880         8192
.data          81920       303104
.bss           11592       385024
Total         388392

size  :
section         size         addr
.text         294880         8192
.data          81920       303104

               .bss           11888       385024
               Total         388688

--help Show a summary of acceptable arguments and options.

-d
-o
-x
--radix=number

Using one of these options, you can control whether the size of each section is given in decimal (‘-d’, or ‘--radix=10’); octal (‘-o’, or ‘--radix=8’); or hexadecimal (‘-x’, or ‘--radix=16’). In ‘--radix=number’, only the three values (8, 10, 16) are supported. The total size is always given in two radices; decimal and hexadecimal for ‘-d’ or ‘-x’ output, or octal and hexadecimal if you’re using ‘-o’.

--target=bfdname

Specify that the object-code format for objfile is bfdname. This option may not be necessary; size can automatically recognize many formats. See Chapter 6 [objdump], page 15, for information on listing available formats.

-V
--version

Display the version number of size.

9 strip

     strip [ -F format | --format=format | --target=format ]
           [ -I format | --input-format=format ]
           [ -O format | --output-format=format ]
           [ -s | --strip-all ] [ -S | -g | --strip-debug ]
           [ -x | --discard-all ] [ -X | --discard-locals ]
           [ -v | --verbose ]  [ -V | --version ]

objfile…

GNU strip discards all symbols from object files objfile. The list of object files may include archives.

strip will not execute unless at least one object file is listed.

strip modifies the files named in its argument, rather than writing modified copies under different names.

-I format

--input-format=format

Treat the original objfile as a file with the object code format format.

-O format

--output-format=format

Replace objfile with a file in the output format format.

-F format

--format=format
--target=format

Treat the original objfile as a file with the object code format format, and rewrite it in the same format.

-s
--strip-all

Remove all symbols.

-g
-S
--strip-debug

Remove debugging symbols only.

-x
--discard-all

Remove non-global symbols.

-X
--discard-locals

Remove compiler-generated local symbols. (These usually start with ‘L’ or ‘.’.)

-V
--version

Show the version number for strip.

-v
--verbose

Verbose output: list all object files modified. In the case of archives, ‘strip -v’ lists all members of the archive.

10 c++filt

The C++ language provides function overloading, which means that you can write many function with the same name (but taking different kinds of parameters). So that the linker can keep these overloaded functions from clashing, all C++ function names are encoded (“mangled”) into a funnylooking low-level assembly label. The c++filt program does the inverse mapping: It decodes (“demangles”) low-level names into user-level names.

When you use c++filt as a filter (which is usually the case), it reads from standard input. Every alphanumeric word (consisting of letters, digits, underscores, dollars, or periods) seen in the input is a potential label. If the label decodes into a C++ name. the C++ name will replace the low-level name in the output.

A typical use of c++filt is to pipe the output of nm though it.

Note that on some systems, both the C and C++ compilers put an underscore in front of every name. (I.e. the C name foo gets the low-level name _foo.) On such systems, c++filt removes any initial underscore of a potential label.

24 GNU Binary Utilities Index 25

Index

.

.stab . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

A

a.out . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 all header information, object file . . . . . . . . . . . . . . . . . . 16 ar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 ar compatibility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 architecture. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 architectures available . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 archive contents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 archive headers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 archives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

C

c++filt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 collections of files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 compatibility, ar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 contents of archive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 creating archives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

D

dates in archive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 debug symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 debugging symbols. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 deleting from archive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 demangling C++ symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 disassembling object code . . . . . . . . . . . . . . . . . . . . . . . . . . 15 discarding symbols. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

E

ELF object file format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 external symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13, 14 extract from archive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

F

file name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

H

header information, all . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

I

input file name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

L

ld . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 libraries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 linker . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

M

machine instructions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 moving in archive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 MRI compatibility, ar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

N

name duplication in archive . . . . . . . . . . . . . . . . . . . . . . . . . 5 name length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 nm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

O

objdump . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 object code format . . . . . . . . . . . . . . . . . . . . . . . . . 14, 15, 20 object file header . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 object file information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 object file sections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 object formats available . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 operations on archive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

P

printing from archive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

Q

quick append to archive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

R

radix for section sizes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 ranlib. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

26 GNU Binary Utilities

relative placement in archive . . . . . . . . . . . . . . . . . . . . . . . . 5 relocation entries, in object file. . . . . . . . . . . . . . . . . . . . . 16 removing symbols. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 repeated names in archive . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 replacement in archive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

S

scripts, ar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 section headers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 section information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 section sizes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 sections, full contents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 size display format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 size number format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 sorting symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 source file name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 source filenames for object files. . . . . . . . . . . . . . . . . . . . . 16 stab . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 strip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 symbol index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3, 17 symbol index, listing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 symbol table entries, printing . . . . . . . . . . . . . . . . . . . . . . 16 symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 symbols, discarding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

U

undefined symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 Unix compatibility, ar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 updating an archive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

V

version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

W

writing archive index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

Table of Contents

1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 ar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

2.1 Controlling ar on the command line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.2 Controlling ar with a script . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

3 objcopy. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 4 ld . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 5 nm. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 6 objdump . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 7 ranlib. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 8 size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 9 strip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 10 c++filt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

ii GNU Binary Utilities